OpenStudy (anonymous):

Find the product: (x + 2)( x − 2)

OpenStudy (akshay_budhkar):

\[(a+b)(a-b)= a^2-b^2\]

OpenStudy (anonymous):

I just figured I would do this in detail (x + 2)(x - 2) x * x= x^2 x * -2= -2x 2* x= 2x 2* -2= -4 x^2 - 2x + 2x -4 (the 2x are cancelled out) x^2 -4